# Majorana fermion induced power-law scaling in the violations of the Wiedemann-Franz law

### v2

@@ IV.2 Inelastic scattering with only phase relaxation
${\begin{pmatrix}I_{1c}^{\downarrow,e}\\ I_{1c}^{\downarrow,h}\\ I_{2c}^{\downarrow,e}\\ I_{2c}^{\downarrow,h}\\ I_{3c}^{\downarrow,e}\\ I_{3c}^{\downarrow,h}\\ I_{4c}^{\downarrow,e}\\ I_{4c}^{\downarrow,h}\end{pmatrix}=-\begin{pmatrix}G(\mathcal{T}^{\downarrow;e}_{12})&G(\mathcal{T}^{\downarrow;e}_{13})&G(\mathcal{T}^{\downarrow;e}_{14})&S(\mathcal{T}^{\downarrow;e}_{12})&S(\mathcal{T}^{\downarrow;e}_{13})&S(\mathcal{T}^{\downarrow;e}_{14})\\ G(\mathcal{T}^{\downarrow;h}_{12})&G(\mathcal{T}^{\downarrow;h}_{13})&G(\mathcal{T}^{\downarrow;h}_{14})&S(\mathcal{T}^{\downarrow;h}_{12})&S(\mathcal{T}^{\downarrow;h}_{13})&S(\mathcal{T}^{\downarrow;h}_{14})\\ G(\mathcal{T}^{\downarrow;e}_{22})&G(\mathcal{T}^{\downarrow;e}_{23})&G(\mathcal{T}^{\downarrow;e}_{24})&S(\mathcal{T}^{\downarrow;e}_{22})&S(\mathcal{T}^{\downarrow;e}_{23})&S(\mathcal{T}^{\downarrow;e}_{24})\\ G(\mathcal{T}^{\downarrow;h}_{22})&G(\mathcal{T}^{\downarrow;h}_{23})&G(\mathcal{T}^{\downarrow;h}_{24})&S(\mathcal{T}^{\downarrow;h}_{22})&S(\mathcal{T}^{\downarrow;h}_{23})&S(\mathcal{T}^{\downarrow;h}_{24})\\ G(\mathcal{T}^{\downarrow;e}_{32})&G(\mathcal{T}^{\downarrow;e}_{33})&G(\mathcal{T}^{\downarrow;e}_{34})&S(\mathcal{T}^{\downarrow;e}_{32})&S(\mathcal{T}^{\downarrow;e}_{33})&S(\mathcal{T}^{\downarrow;e}_{34})\\ G(\mathcal{T}^{\downarrow;h}_{32})&G(\mathcal{T}^{\downarrow;h}_{33})&G(\mathcal{T}^{\downarrow;h}_{34})&S(\mathcal{T}^{\downarrow;he}_{32})&S(\mathcal{T}^{\downarrow;h}_{33})&S(\mathcal{T}^{\downarrow;h}_{34})\\ G(\mathcal{T}^{\downarrow;e}_{42})&G(\mathcal{T}^{\downarrow;e}_{43})&G(\mathcal{T}^{\downarrow;e}_{44})&S(\mathcal{T}^{\downarrow;e}_{42})&S(\mathcal{T}^{\downarrow;e}_{43})&S(\mathcal{T}^{\downarrow;e}_{44})\\ G(\mathcal{T}^{\downarrow;h}_{42})&G(\mathcal{T}^{\downarrow;h}_{43})&G(\mathcal{T}^{\downarrow;h}_{44})&S(\mathcal{T}^{\downarrow;h}_{42})&S(\mathcal{T}^{\downarrow;h}_{43})&S(\mathcal{T}^{\downarrow;h}_{44})\end{pmatrix}\begin{pmatrix}V\\ V_{3}\\ V_{3}\\ \Delta_{R}\\ \Delta T\\ \Delta T\end{pmatrix}},$ (42)
and the corresponding heat currents due to spin-down electrons and holes, as,
${\begin{pmatrix}I_{1q}^{\downarrow,e}\\ I_{1q}^{\downarrow,h}\\ I_{2q}^{\downarrow,e}\\ I_{2q}^{\downarrow,h}\\ I_{3q}^{\downarrow,e}\\ I_{3q}^{\downarrow,h}\\ I_{4q}^{\downarrow,e}\\ I_{4q}^{\downarrow,h}\end{pmatrix}=-\begin{pmatrix}TS(\mathcal{T}^{\downarrow;e}_{12})&TS(\mathcal{T}^{\downarrow;e}_{13})&TS(\mathcal{T}^{\downarrow;e}_{14})&L(\mathcal{T}^{\downarrow;e}_{12})&L(\mathcal{T}^{\downarrow;e}_{13})&L(\mathcal{T}^{\downarrow;e}_{14})\\ TS(\mathcal{T}^{\downarrow;h}_{12})&TS(\mathcal{T}^{\downarrow;h}_{13})&TS(\mathcal{T}^{\downarrow;h}_{14})&L(\mathcal{T}^{\downarrow;h}_{12})&L(\mathcal{T}^{\downarrow;h}_{13})&L(\mathcal{T}^{\downarrow;h}_{14})\\ TS(\mathcal{T}^{\downarrow;e}_{22})&TS(\mathcal{T}^{\downarrow;e}_{23})&TS(\mathcal{T}^{\downarrow;e}_{24})&L(\mathcal{T}^{\downarrow;e}_{22})&L(\mathcal{T}^{\downarrow;e}_{23})&L(\mathcal{T}^{\downarrow;e}_{24})\\ TS(\mathcal{T}^{\downarrow;h}_{22})&TS(\mathcal{T}^{\downarrow;h}_{23})&TS(\mathcal{T}^{\downarrow;h}_{24})&L(\mathcal{T}^{\downarrow;h}_{22})&L(\mathcal{T}^{\downarrow;h}_{23})&L(\mathcal{T}^{\downarrow;h}_{24})\\ TS(\mathcal{T}^{\downarrow;e}_{32})&TS(\mathcal{T}^{\downarrow;e}_{33})&TS(\mathcal{T}^{\downarrow;e}_{34})&L(\mathcal{T}^{\downarrow;e}_{32})&L(\mathcal{T}^{\downarrow;e}_{33})&L(\mathcal{T}^{\downarrow;e}_{34})\\ TS(\mathcal{T}^{\downarrow;h}_{32})&TS(\mathcal{T}^{\downarrow;h}_{33})&TS(\mathcal{T}^{\downarrow;h}_{34})&L(\mathcal{T}^{\downarrow;he}_{32})&L(\mathcal{T}^{\downarrow;h}_{33})&L(\mathcal{T}^{\downarrow;h}_{34})\\ TS(\mathcal{T}^{\downarrow;e}_{42})&TS(\mathcal{T}^{\downarrow;e}_{43})&TS(\mathcal{T}^{\downarrow;e}_{44})&L(\mathcal{T}^{\downarrow;e}_{42})&L(\mathcal{T}^{\downarrow;e}_{43})&L(\mathcal{T}^{\downarrow;e}_{44})\\ TS(\mathcal{T}^{\downarrow;h}_{42})&TS(\mathcal{T}^{\downarrow;h}_{43})&TS(\mathcal{T}^{\downarrow;h}_{44})&L(\mathcal{T}^{\downarrow;h}_{42})&L(\mathcal{T}^{\downarrow;h}_{43})&L(\mathcal{T}^{\downarrow;h}_{44})\end{pmatrix}\begin{pmatrix}V\\ V_{3}\\ V_{3}\\ \Delta_{R}\\ \Delta T\\ \Delta T\end{pmatrix}},$ (43)
where $G(X)$ is the electrical conductance for particles with transmission probability $X$, $S(X)$ is the Seebeck coefficient for particles with transmission probability $X$, and $L(X)$ is the thermal conductance for particles with transmission probability $X$. The electrical, Seebeck, and thermal conductances are given in Eqs. (21-23). Similar to the derivation of the Onsager matrix elements for the inelastic scatterer with both phase and momentum relaxation, one can find the Onsager matrix elements for the inelastic scatterer with only phase relaxation. First, we write Eqs. (40-43) in terms of $V$, and $\Delta T_{R}$ only by using the condition of the Buttiker voltage-temperature probe ($I_{3c}(V,T+\Delta T)=I_{3q}(V,T+\Delta T)=I_{4c}(V,T+\Delta T)=I_{4q}(V,T+\Delta T)=0$). This allows us to write $I_{2c}$, and $I_{2q}$ in terms of $V$ and $\Delta T$ only as,
$\begin{split}\small I_{2c}=L^{\prime\prime}_{cV}V+L^{\prime\prime}_{cT}\Delta T_{R}\text{, and }I_{2q}=L^{\prime\prime}_{qV}V+L^{\prime\prime}_{qT}\Delta T_{R}.\\ \end{split}$ (44)
where $\Delta=\sum_{s,k}\bigg{[}(G_{33}^{s,k}+G_{34}^{s,k}+G_{43}^{s,k}+G_{44}^{s,k})(L_{33}^{s,k}+L_{34}^{s,k}+L_{43}^{s,k}+L_{44}^{s,k})-(S_{33}^{s,k}+S_{34}^{s,k}+S_{43}^{s,k}+S_{44}^{s,k})^{2}\bigg{]}$. The effective Onsager matrix elements can find all the thermoelectric coefficients. For the setup with the Buttiker voltage-temperature probe with phase relaxation only, the thermal conductance is given by [30, 31],
The Lorenz ratio for the setup with the Buttiker voltage-temperature probe with phase relaxation only is then given by,
The Mathematica codes are available in GitHub  [32].
Figure 2: The Lorentz ratio ($W$) vs. the coupling strength of the inelastic scatterer $\epsilon$ for (a) Buttiker 3 $\times$ 3 scatterer with both phase and momentum relaxation (b) Buttiker 4 $\times$ 4 scatterer with only phase relaxation
Setup Coupled MBS Individual MBS
Topological Kondo model [6] N/A $W=\frac{0.66}{g}$
MBS ABI with both phase and momentum relaxation
Buttiker $3\times 3$ scatterer [8]
$W\approx\frac{0.74}{\epsilon^{1/5}}$ $W\approx\frac{0.76}{\epsilon^{1/4}}$
MBS ABI with only phase relaxation
Buttiker $4\times 4$ scatterer [9]
$W\approx\frac{0.7}{\epsilon^{2/9}}$ $W\approx\frac{0.6}{\epsilon^{2/7}}$
V Analysis: Violation of WF law phenomenologically vs. Luttinger liquid model
In Fig. 2, we plot the Lorenz ratio vs. the strength of inelastic scattering for the setup with both phase and momentum relaxation [8], and the setup with only phase relaxation [9] in the presence of coupled MBS, individual MBS, and the absence of MBS. We use the parameters, Fermi energy $E_{F}=10\mu eV$, and flux $\phi=\phi_{0}$, $\phi_{0}$ being the flux quantum $hc/e$. In the absence of MBS, WF law is preserved in all three cases, i.e., $W=1$ for the $3\times 3$ Buttiker scatterer with both phase and momentum relaxation (Fig. 2 (a)), and the $4\times 4$ Buttiker scatterer with only phase relaxation (Fig. 2 (b)) regardless of inelastic scattering $\epsilon$. For the presence of MBS, we consider two cases: individual MBS and coupled MBS. In our setup, we consider the MBS coupling strength $E_{M}=10\mu eV$ when considering coupled MBS. The violation in Wiedemann-Franz law is highest when $E_{F}$ is close to $E_{M}$. Thus, to distinguish between individual MBS ($E_{M}=0$), and coupled MBS ($E_{M}=10\mu eV$), we take $E_{F}=10\mu eV$.
For the $3\times 3$ Buttiker S-matrix [8], and $4\times 4$ Buttiker S-matrix [9], for $\epsilon\to 0$, the Buttiker voltage-temperature probe is completely disconnected, and there is no inelastic scattering in the setup. When MBS are uncoupled or individual ($E_{M}=0$), the Lorenz ratio ($W$) is one at the limit of no inelastic scattering and decays with increasing strength of inelastic scattering ($\epsilon$). For the Luttinger liquid setup with individual MBS studied in Ref. [6], the Lorenz ratio $W$ passes through one at $g=1$, i.e., in the absence of inelastic scattering. In our setup, the Lorenz ratio $W$ is one in the absence of inelastic scattering for individual MBS. Thus, our results for individual MBS corroborate the findings of the Luttinger liquid model studied in Ref. [6]. For the $3\times 3$ Buttiker S-matrix [8] with both phase and momentum relaxation, the Lorenz ratio scales with inelastic scattering ($\epsilon$) as $W\approx\frac{0.76}{\epsilon^{1/4}}$. Finally, for $4\times 4$ Buttiker S-matrix with phase relaxation only [9], the Lorenz ratio scales with $\epsilon$ as $W\approx\frac{0.6}{\epsilon^{2/7}}$. Thus, for individual Majorana, the Lorenz ratio $W$ is inversely proportional to the strength of inelastic scattering $\epsilon$ for inelastic scattering with both phase and momentum relaxation with exponent $1/4$. In contrast, the Lorenz ratio scales inversely with $\epsilon^{2/7}$ for inelastic scattering with phase relaxation only.
For coupled MBS, $E_{M}=10\mu eV\neq 0$, we see that the Lorenz ratio is greater than one in the absence of inelastic scattering for all three cases, i.e. $W>1$ for $\epsilon\to 0$ for the $3\times 3$ Buttiker S-matrix [8] with both phase and momentum relaxation, and $\epsilon\to 0$ for the $4\times 4$ Buttiker S-matrix with only phase relaxation. In the presence of coupled MBS, PHS is broken such that the majority of electrons and holes travel in the same direction. It leads to a reduction in the net charge current and a commensurate increase in the heat current. Thus, in the presence of coupled MBS, WF law is violated even in the limit of zero inelastic scattering. The Lorenz ratio decays with increasing $\epsilon$ universally in the presence of MBS, regardless of whether they are coupled or uncoupled. For the $3\times 3$ Buttiker S-matrix with both phase and momentum relaxation [8], the Lorenz ratio passes through one at $\epsilon=0.13$ (see Fig. 2 (a)). While for the $4\times 4$ Buttiker S-matrix with phase relaxation only, the Lorenz ratio passes through one at $\epsilon=0.08$ (see Fig. 2 (b)). Thus, in the presence of coupled MBS, inelastic scattering can recover the WF law at a particular $\epsilon$. From Fig. 2, we find that the Lorenz ratio $W$ scaling with respect to inelastic scattering $\epsilon$ also follows the power-law for coupled MBS. For the $3\times 3$ Buttiker S-matrix [8] with both phase and momentum relaxation (Fig. 2 (a)), the Lorenz ratio scales with inelastic scattering ($\epsilon$) as $W\approx\frac{0.74}{\epsilon^{1/5}}$. For $4\times 4$ Buttiker S-matrix with phase relaxation only [9], the Lorenz ratio scales with $\epsilon$ as $W\approx\frac{0.7}{\epsilon^{2/9}}$. Thus, the Lorenz ratio for coupled MBS is inversely proportional to $\epsilon^{1/5}$ when both phase and momentum relaxation are present for the Buttiker scatterer. In the presence of inelastic scattering with phase relaxation only, the Lorenz ratio is inversely proportional to $\epsilon^{2/9}$. This not only shows a way to detect edge modes. It confirms that coupled and individual MBS have different nature as function of inelastic scattering and scaling of the violations are distinct..
We compare our results with the Luttinger liquid model studied in Ref. [6] in Table I. The setup in Ref. [6] considers a quantum dot junction hosting MBS in the topological Kondo regime using the many-body formalism. The Luttinger parameter $g$ describes electron-electron interaction in the many-body picture with $g=1$ corresponding to no interaction. The authors in Ref. [6] report that weakly coupled MBS ($E_{M}\approx 0$) violate WF law. The authors show that the Lorenz ratio scales as $W\approx\frac{2}{3g}$, i.e., the Lorenz ratio is inversely proportional to the Luttinger parameter in the presence of uncoupled MBS. In contrast in our setup, the Lorentz ratio is inversely proportional to $\epsilon^{1/4}$ for Buttiker scatterer when both phase and momentum relaxation are present. Similarly, for the phase relaxation case, it is inversely proportional to $\epsilon^{2/7}$. Thus, Table I shows the power law valid for our setup, either with phase and momentum relaxation or with only phase relaxation. In either case, one gets distinct power law violations for uncoupled and coupled MBS. This provides another unique method to detect MBS in mesoscopic superconducting junctions.
VI Experimental realization & Conclusion
[…]
To conclude, we studied the violation of Wiedemann-Franz law in the presence and absence of MBS and the scaling of the Lorenz ratio concerning the strength of inelastic scattering induced by a Buttiker voltage-temperature probe. We find that WF law is only violated in the presence of MBS. We studied the scaling of the Lorenz ratio in the presence of inelastic scattering with both phase and momentum relaxation and with only phase relaxation and compared our results with those of the Luttinger liquid model studied in Ref. [6]. We showed that the Lorenz ratio decays with increasing strength of inelastic scattering when MBS are present, irrespective of whether they are individual or coupled. The Luttinger liquid model predicts that the Lorenz ratio is inversely proportional to the Luttinger parameter with $W(g)\approx\frac{2}{3g}$. In the presence of individual MBS, for the $3\times 3$ Buttiker S-matrix [8] with both phase and momentum relaxation, the Lorenz ratio scales with inelastic scattering ($\epsilon$) as $W\approx\frac{0.76}{\epsilon^{1/4}}$. For $4\times 4$ Buttiker S-matrix with phase relaxation only [9], the Lorenz ratio scales with $\epsilon$ as $W\approx\frac{0.6}{\epsilon^{2/7}}$. When MBS in the setup are coupled, the scaling is similar with $W\approx\frac{0.74}{\epsilon^{1/5}}$ for the $3\times 3$ Buttiker S-matrix [8], and $W\approx\frac{0.7}{\epsilon^{2/9}}$ for $4\times 4$ Buttiker S-matrix with phase relaxation only [9]. The results show that the electron-electron interaction in the Luttinger liquid model is similar to the inelastic scattering induced phenomenologically by the Buttiker voltage-temperature probe [10] with both phase and momentum relaxation [8, 13]. Further, we show that for coupled MBS, the Lorenz ratio is greater than one when inelastic scattering is absent and decays with increasing $\epsilon$. For individual MBS, the Lorenz ratio is conserved without inelastic scattering and decays when inelastic scattering is introduced. This distinct behavior of the Lorenz ratio in the scaling can be used to detect MBS.

### v4

@@ Appendix B Expression for charge and heat current and Onsager coefficients with phase relaxation
${\begin{pmatrix}I_{1c}^{\downarrow,e}\\ I_{1c}^{\downarrow,h}\\ I_{2c}^{\downarrow,e}\\ I_{2c}^{\downarrow,h}\\ I_{3c}^{\downarrow,e}\\ I_{3c}^{\downarrow,h}\\ I_{4c}^{\downarrow,e}\\ I_{4c}^{\downarrow,h}\end{pmatrix}=-\begin{pmatrix}G(\mathcal{T}^{\downarrow;e}_{12})&G(\mathcal{T}^{\downarrow;e}_{13})&G(\mathcal{T}^{\downarrow;e}_{14})&S(\mathcal{T}^{\downarrow;e}_{12})&S(\mathcal{T}^{\downarrow;e}_{13})&S(\mathcal{T}^{\downarrow;e}_{14})\\ G(\mathcal{T}^{\downarrow;h}_{12})&G(\mathcal{T}^{\downarrow;h}_{13})&G(\mathcal{T}^{\downarrow;h}_{14})&S(\mathcal{T}^{\downarrow;h}_{12})&S(\mathcal{T}^{\downarrow;h}_{13})&S(\mathcal{T}^{\downarrow;h}_{14})\\ G(\mathcal{T}^{\downarrow;e}_{22})&G(\mathcal{T}^{\downarrow;e}_{23})&G(\mathcal{T}^{\downarrow;e}_{24})&S(\mathcal{T}^{\downarrow;e}_{22})&S(\mathcal{T}^{\downarrow;e}_{23})&S(\mathcal{T}^{\downarrow;e}_{24})\\ G(\mathcal{T}^{\downarrow;h}_{22})&G(\mathcal{T}^{\downarrow;h}_{23})&G(\mathcal{T}^{\downarrow;h}_{24})&S(\mathcal{T}^{\downarrow;h}_{22})&S(\mathcal{T}^{\downarrow;h}_{23})&S(\mathcal{T}^{\downarrow;h}_{24})\\ G(\mathcal{T}^{\downarrow;e}_{32})&G(\mathcal{T}^{\downarrow;e}_{33})&G(\mathcal{T}^{\downarrow;e}_{34})&S(\mathcal{T}^{\downarrow;e}_{32})&S(\mathcal{T}^{\downarrow;e}_{33})&S(\mathcal{T}^{\downarrow;e}_{34})\\ G(\mathcal{T}^{\downarrow;h}_{32})&G(\mathcal{T}^{\downarrow;h}_{33})&G(\mathcal{T}^{\downarrow;h}_{34})&S(\mathcal{T}^{\downarrow;he}_{32})&S(\mathcal{T}^{\downarrow;h}_{33})&S(\mathcal{T}^{\downarrow;h}_{34})\\ G(\mathcal{T}^{\downarrow;e}_{42})&G(\mathcal{T}^{\downarrow;e}_{43})&G(\mathcal{T}^{\downarrow;e}_{44})&S(\mathcal{T}^{\downarrow;e}_{42})&S(\mathcal{T}^{\downarrow;e}_{43})&S(\mathcal{T}^{\downarrow;e}_{44})\\ G(\mathcal{T}^{\downarrow;h}_{42})&G(\mathcal{T}^{\downarrow;h}_{43})&G(\mathcal{T}^{\downarrow;h}_{44})&S(\mathcal{T}^{\downarrow;h}_{42})&S(\mathcal{T}^{\downarrow;h}_{43})&S(\mathcal{T}^{\downarrow;h}_{44})\end{pmatrix}\begin{pmatrix}V\\ V_{3}\\ V_{3}\\ \Delta T_{R}\\ \Delta T\\ \Delta T\end{pmatrix}},$ (49)
and the corresponding heat currents due to spin-down electrons and holes as,
${\begin{pmatrix}I_{1q}^{\downarrow,e}\\ I_{1q}^{\downarrow,h}\\ I_{2q}^{\downarrow,e}\\ I_{2q}^{\downarrow,h}\\ I_{3q}^{\downarrow,e}\\ I_{3q}^{\downarrow,h}\\ I_{4q}^{\downarrow,e}\\ I_{4q}^{\downarrow,h}\end{pmatrix}=-\begin{pmatrix}TS(\mathcal{T}^{\downarrow;e}_{12})&TS(\mathcal{T}^{\downarrow;e}_{13})&TS(\mathcal{T}^{\downarrow;e}_{14})&L(\mathcal{T}^{\downarrow;e}_{12})&L(\mathcal{T}^{\downarrow;e}_{13})&L(\mathcal{T}^{\downarrow;e}_{14})\\ TS(\mathcal{T}^{\downarrow;h}_{12})&TS(\mathcal{T}^{\downarrow;h}_{13})&TS(\mathcal{T}^{\downarrow;h}_{14})&L(\mathcal{T}^{\downarrow;h}_{12})&L(\mathcal{T}^{\downarrow;h}_{13})&L(\mathcal{T}^{\downarrow;h}_{14})\\ TS(\mathcal{T}^{\downarrow;e}_{22})&TS(\mathcal{T}^{\downarrow;e}_{23})&TS(\mathcal{T}^{\downarrow;e}_{24})&L(\mathcal{T}^{\downarrow;e}_{22})&L(\mathcal{T}^{\downarrow;e}_{23})&L(\mathcal{T}^{\downarrow;e}_{24})\\ TS(\mathcal{T}^{\downarrow;h}_{22})&TS(\mathcal{T}^{\downarrow;h}_{23})&TS(\mathcal{T}^{\downarrow;h}_{24})&L(\mathcal{T}^{\downarrow;h}_{22})&L(\mathcal{T}^{\downarrow;h}_{23})&L(\mathcal{T}^{\downarrow;h}_{24})\\ TS(\mathcal{T}^{\downarrow;e}_{32})&TS(\mathcal{T}^{\downarrow;e}_{33})&TS(\mathcal{T}^{\downarrow;e}_{34})&L(\mathcal{T}^{\downarrow;e}_{32})&L(\mathcal{T}^{\downarrow;e}_{33})&L(\mathcal{T}^{\downarrow;e}_{34})\\ TS(\mathcal{T}^{\downarrow;h}_{32})&TS(\mathcal{T}^{\downarrow;h}_{33})&TS(\mathcal{T}^{\downarrow;h}_{34})&L(\mathcal{T}^{\downarrow;he}_{32})&L(\mathcal{T}^{\downarrow;h}_{33})&L(\mathcal{T}^{\downarrow;h}_{34})\\ TS(\mathcal{T}^{\downarrow;e}_{42})&TS(\mathcal{T}^{\downarrow;e}_{43})&TS(\mathcal{T}^{\downarrow;e}_{44})&L(\mathcal{T}^{\downarrow;e}_{42})&L(\mathcal{T}^{\downarrow;e}_{43})&L(\mathcal{T}^{\downarrow;e}_{44})\\ TS(\mathcal{T}^{\downarrow;h}_{42})&TS(\mathcal{T}^{\downarrow;h}_{43})&TS(\mathcal{T}^{\downarrow;h}_{44})&L(\mathcal{T}^{\downarrow;h}_{42})&L(\mathcal{T}^{\downarrow;h}_{43})&L(\mathcal{T}^{\downarrow;h}_{44})\end{pmatrix}\begin{pmatrix}V\\ V_{3}\\ V_{3}\\ \Delta T_{R}\\ \Delta T\\ \Delta T\end{pmatrix}},$ (50)
where $G(X)$ is the electrical conductance for particles with transmission probability $X$, $S(X)$ is the Seebeck coefficient for particles with transmission probability $X$, and $L(X)$ is the thermal conductance for particles with transmission probability $X$. The electrical, Seebeck, and thermal conductances are given in Eqs. (40-42).